If one card is drawn from an ordinary deck of cards, find the probability of getting
each event:
(a) A 7 or an 8 or a 9
(b) A spade or a queen or a king
(c) A club or a face card
(d) A 9 or a 10 or a spade or a club

Answer by greenestamps(13200)



(a) A 7 or an 8 or a 9

4 7's, 4 8's, and 4 9's, so a total of 12. P=12/52 = 3/13

(b) A spade or a queen or a king

13 spades; 3 queens that are not spades; 3 kings that are not spades -- total 19. P=19/52

(c) A club or a face card

13 clubs; 3*3=9 face cards that are not clubs -- total 22. P=22/52 = 11/26

(d) A 9 or a 10 or a spade or a club

4 9's and 4 10's; 11 spades that are neither 9 or 10; 11 clubs that are neither 9 or 10 -- total 30. P=30/52 = 15/26

OR...

13 spades; 13 clubs; 2 9's that are neither spade not club; 2 10's that are neither spade not club -- total 30. So again P=30/52 = 15/26.